# ASGO: Adaptive Structured Gradient Optimization

### v1

@@ Abstract
Training deep neural networks (DNNs) is a structured optimization problem, because the parameters are naturally represented by matrices and tensors rather than simple vectors. Under this structural representation, it has been widely observed that gradients are low-rank and Hessians are approximately block-wise diagonal. These structured properties are crucial for designing efficient optimization algorithms but may not be utilized by current popular optimizers like Adam. In this paper, we present a novel optimization algorithm ASGO that capitalizes on these properties by employing a preconditioner that is adaptively updated using structured gradients. By fine-grained theoretical analysis, ASGO is proven to achieve superior convergence rates compared to existing structured gradient methods. Based on the convergence theory, we further demonstrate that ASGO can benefit from the low-rank and block-wise diagonal properties. We also discuss practical modifications of ASGO and empirically verify the effectiveness of the algorithm on language model tasks.
@@ 1 Introduction
Numerical optimization algorithms, especially those can efficiently train large foundation models (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023, ; Ouyang et al., 2022), play an important role in the modern machine learning field. Among them, adaptive gradient methods like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2014) are popular choices, gaining huge success in training state-of-the-art models in many tasks. These algorithms typically uses a diagonal matrix preconditioner in the update, like for AdaGrad
@@
This coordinate-wise step size design has been theoretically verified to be effective as it can exploit the sparsity of the gradient vectors $g_{t}$ (Duchi et al., 2011). Also, when the Hessian is well-approximated by a diagonal matrix whose diagonal entries have very different scales, adaptive gradient methods have been proven to be beneficial (Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). While these results seem to be convincing, common deep neural networks (DNNs) do not necessarily have sparse gradients or Hessians that are well-approximated by ill-conditioned diagonal matrices. Instead, if we take the matrix structure of gradients in neural networks into account, it has been widely observed that these structured gradients are usually low-rank (Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023), and the Hessians are well-approximated by block-wise diagonal matrices (Collobert, 2004; Zhang et al., 2024a, b). Since adaptive gradient methods simply treat the parameters as vectors and ignore the matrix structure of gradients, they are generally unable to exploit these structured properties. This gap makes us ask the following question: How can we properly consider the matrix structures of gradients and exploit the low-rank and block-wise diagonal properties?
One possible attempt is Shampoo (Gupta et al., 2018), which considers the matrix (or tensor) structure of gradients:
$W_{t+1}=W_{t}-\eta_{t}L_{t}^{-\frac{1}{4}}G_{t}R_{t}^{-\frac{1}{4}},\text{ where $W_{t}\in\mathbb{R}^{m\times n}$ and $L_{t}\in\mathbb{R}^{m\times m}$, $R_{t}\in\mathbb{R}^{n\times n}$ are full matrices.}$.
@@
The main motivation for such design is that if we apply vectorization for the update, the Shampoo preconditioner is a single matrix that is the Kronecker product of $L_{t}$ and $R_{t}$, which is an approximation of the full-matrix preconditioner for AdaGrad (Duchi et al., 2011). However, the theoretical convergence of Shampoo is worse than that for AdaGrad or even SGD when the dimension is large. Also, Shampoo needs more memory and much heavier computation than adaptive gradient methods because it requires two preconditioners, making it less suitable for training large-scale DNNs.
@@
In this paper, we provide an answer to the aforementioned question by proposing ASGO (Adaptive Structured Gradient Optimization), which significantly improves the convergence guarantees of Shampoo while requiring less memory and computation. In light of the analysis that demonstrates the benefits of adaptive gradient methods (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2020), we use appropriate assumptions to enable a more fine-grained convergence analysis, showing superior convergence results and how ASGO can benefit from the low-rank and block-wise diagonal properties of the problem. We also discuss the connection between ASGO and Muon (Jordan et al., 2024), a structured gradient method originated from the steepest descent algorithm with spectral norm, to conjecture a relation between ASGO and Muon analogous to the relation between AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). Furthermore, we develop an efficient design for query-key attention parameters in transformer models and examine the empirical performance on pretraining transformer model tasks.
Our main contributions are summarized as follows:
@@
  • We propose ASGO, an algorithm utilizing the structured gradient properties. We theoretically analyze the convergence of ASGO and show that it converges faster than full-matrix AdaGrad and Shampoo.
@@
  • We further demonstrate that ASGO can effectively exploit the low-rankness of gradients as well as the approximate block-wise diagonal property of Hessians that is typically observed in training DNNS, hence indicating that ASGO has a great potential for real-world applications.
@@
  • We develop a practical implementation of ASGO with targeted modifications for transformer architectures that offers significant advantages: it eliminates the need for a separate optimizer for 1D parameters like Muon, while requiring less memory and computational complexity compared to Shampoo.
@@
  • We empirically validate the effectiveness of ASGO and the modifications on Transformer pretraining tasks, demonstrating the algorithm’s great potential in deep learning tasks.
@@ 2 Related Work
In the standard optimization literature (i.e., excluding areas such as conic optimization), it is common to consider variables as vectors. However, recently, optimization methods for machine learning that consider variables as matrices have been rapidly gaining attention. Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) consider the matrix or layer structure to help reduce the memory cost of Adam and enable more efficient training. KFAC (Martens and Grosse, 2015) and Shampoo (Gupta et al., 2018) are two pioneering works that approximate, respectively, the full-matrix preconditioner of AdaGrad and the empirical Fisher matrix, using Kronecker products of smaller matrices, making the memory cost more affordable. Unfortunately, the analysis in (Gupta et al., 2018) shows that the rate of convergence for Shampoo is no better than that for the full-matrix AdaGrad. We note that there is another method, TNT (Ren and Goldfarb, 2021) that is very closely related to Shampoo. TNT was developed as a natural gradient method, approximating the true Fisher matrix (FM) by the covariance of block-wise sampling-based gradients assuming that they are Tensor-Normally distributed. The main difference between TNT and Shampoo, besides that between the FM and EF matrices, is that TNT uses the inverse of their Kronecker factors, whereas Shampoo uses the -1/2 power of them.
Many recent works have focused on improving full-matrix AdaGrad and Shampoo. Feinberg et al. (2023) uses a sketching-based approach to approximate the full-matrix AdaGrad preconditioner with lower memory cost. Morwani et al. (2024) provides theoretical intuition and empirical evidence to claim that Shampoo should use the $-1/2$ power on the preconditioners to better approximate the EF matrix. Vyas et al. (2024) demonstrates that Shampoo is like doing Adafactor in the eigenspace of gradients and proposes a novel algorithm, SOAP, that performs Adam in this eigenspace. SOAP is observed to achieve better performance than Adam and Shampoo, but suffers from a high computation load per iteration. Galore (Zhao et al., 2024) shares a similar algorithmic design with SOAP with extra focus on lowering memory costs. Muon (Jordan et al., 2024) follows this line of work, using the steepest descent framework with the spectral norm, and is shown to be scalable and effective in training large foundation models (Liu et al., 2025a). More recently, Nguyen et al. (2025) proposes AdaDiag, which may be viewed as SOAP doing SVD without gradient accumulation. Liu et al. (2025b) proposes COSMOS, a combination of SOAP and Muon, that trades off between performance and computational efficiency.
Rank of Gradients and Weight Updates.
It has been widely observed that gradients are naturally low-rank in DNNs, even when a large batch size is employed (Gur-Ari et al., 2018; Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023). This property has been widely utilized for computation and memory efficiency in training (Wang et al., 2018; Cosson et al., 2023; Zhao et al., 2024). On the other hand, the rank of the total weight update $\Delta W=W_{T}-W_{0}$, actually depends a lot on the training method. If we use LoRA (Hu et al., 2022), $\Delta W$ is determined to be low-rank. However, in pretraining or even many complex fine-tuning tasks, LoRA is observed to perform much worse than methods that produce high-rank weight updates like full-parameter training, which is conjectured to be due to the weight update rank (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025).
Block-wise Diagonal Hessian.
It has been observed that the Hessian of a neural network tends to be block-wise diagonal with each block corresponding to a neuron both in experiments and theory for small MLPs (Collobert, 2004). Zhang et al. (2024a, b) recently numerically verified this property in small transformers and Zhang et al. (2024a) further empirically showed that transformers may exhibit heterogeneity between blocks, while CNNs may not.
Concurrent Work.
When we were finishing writing this paper, we noticed that concurrent work (Xie et al., 2025), which also investigates the same algorithm as ASGO, referred to as One-Sided Shampoo in their paper, just appeared on arXiv. Although the theoretical techniques and convergence results for ASGO are very similar in general, there are still some notable differences between the two works. First, the motivations are different. In Xie et al. (2025), the authors develop ASGO from the unified preconditioning method framework AdaptReg (Gupta et al., 2017) and mainly highlight its superior convergence results compared to Shampoo and full-matrix AdaGrad. In our paper, we focus more on theoretically discussing how ASGO can utilize the structured properties of optimization problems including low-rank gradients and block-wise diagonal Hessians, to highlight the potential of ASGO as a practical algorithm for training deep learning models. Second, our empirical results provide evidence that ASGO can perform well in practical tasks, whereas Xie et al. (2025) focuses on convex settings and examines One-Sided Shampoo only on linear regression. Moreover, our implementation includes specialized designs for Transformer architectures (particularly query/key) to improve performance on deep learning training tasks. As a direct byproduct of our main algorithm, we also developed a lightweight diagonal version named DASGO to trade off memory consumption and performance. Third, Xie et al. (2025) focuses primarily on developing a general proof framework applicable to multiple optimizers including One-Sided Shampoo, whereas our work specifically examines the theoretical and practical benefits of ASGO. To conclude, both works contribute to a better understanding of ASGO/One-Sided Shampoo.
3 Preliminaries and the Algorithm
3.1 Notation
In the following analysis, for an arbitrary matrix $W\in\mathbb{R}^{m\times n}$, we denote
@@ 3.1 Notation
We use ,,$\prec$,$\preceq$,$\succ$,$\succeq$ , ⪯ , ≻ , ⪰ to denote the Lowner order (Löwner, 1934) of square matrices, i.e., for matrices $A,B\in\mathbb{R}^{m\times m}$, $A\preceq B$ implies that $B-A$ is positive semidefinite and $A\prec B$ implies that $B-A$ is positive definite. $\succ$ and $\succeq$ are defined accordingly. Throughout this paper, we typically use capital letters such as $W$ to present matrices and $[W]_{i,j}$ to present the $(i,j)$-th entry of $W$.
3.2 Problem Settings
We study the following stochastic optimization problem:
@@ 3.2 Problem Settings
where we can only access to the stochastic gradient oracle $\nabla f(W;\xi)$ at $W$. In our analysis of convergence, we will mainly consider the convex optimization case.
Assumption 1 (Convexity).
$f(\cdot)$ is convex and $W_{*}$ is one of its minimizers.
Although deep neural networks are in general nonconvex globally, convexity may apply locally in some regions. Thus, convex analysis can be helpful for understanding the behavior of algorithms.
@@
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, schedule $\{\eta_{t}\}$ and batch size $M\in\mathbb{N}$ ($\epsilon$ should be small, similar to the $\epsilon$ for Adam or AdaGrad.)
@@
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$triangleright Compute stochastic (sub)gradient
@@
7:     $\Lambda_{t}=V_{t}^{\frac{1}{2}}+\epsilon I_{m}$triangleright Use the square root of the summation to update
@@
3.3 ASGO (Algorithm 1)
We propose ASGO (Adaptive Structured Gradient Optimization) in Algorithm 1, which is an algorithm with a single-side preconditioner. Compared to full-matrix AdaGrad, ASGO preserves and utilizes the matrix structure of $W_{t}$ and $G_{t}$, avoiding the huge memory cost for the storage of its preconditioner. Compared to Shampoo, ASGO employs only single preconditioner from one side, leading to the following main update rule:
$W_{t+1}=W_{t}-\eta_{t}V_{t}^{-\frac{1}{2}}G_{t},\text{ where $W_{t}\in\mathbb{R}^{m\times n}$ and $V_{t}\in\mathbb{R}^{m\times m}$ is a full matrix,}$,
if we ignore $\epsilon$, which is used for numerical stability. Basically, employing only one preconditioner largely reduces the memory cost and also improves the computational efficiency of Shampoo, as we only need to store one preconditioner matrix and only need to compute one matrix square root and inverse. On the other hand, this change biases from the starting point of Shampoo, as the preconditioner is no longer an effective approximation of the full-matrix AdaGrad preconditioner or the empirical Fisher matrix (Gupta et al., 2018; Morwani et al., 2024). However, as we will see in the following sections, this design can actually better exploit the low-rankness of gradients and the block-wise diagonal nature of Hessians, in terms of achieving better convergence rates.
@@ 4 Nonsmooth Theory
If we additionally assume an upper bound for each stochastic (sub)gradient such that $\mathbb{E}\left[G_{t}G_{t}^{\top}\right]\preceq Q^{2}$, where $Q\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix, Theorem 1 also implies
@@
Let us make some comments regarding the bound. First, one can easily check that the convergence rate should be in general $\mathcal{O}(1/\sqrt{T})$, which is standard for convex nonsmooth problems. An issue here might be that we treat $D_{\mathrm{op}}$ as a constant here, but it may increase as the iteration number $T$ increases. This can be addressed, for example, by involving a projection onto a bounded convex set $\mathcal{W}$, which should be with respect to the norm $\left\|\cdot\right\|_{\Lambda_{t}}$ in each iteration, just like AdaGrad does (Duchi et al., 2011). However, since this projection step is rarely used in training DNNs, we follow Gupta et al. (2018) to omit it in the algorithm to make it more practical. More importantly, this theoretical bound depends on the trace norm of gradients and the spectral norm of weights, showing that the algorithm can make use of the low-rank property of gradients. We would like to share some more points and intuitions about ASGO and the convergence theorem through the following discussions.
  • Comparison with SGD. SGD has the convergence rate of (Zinkevich, 2003):
@@
    where $D_{\mathrm{F}}$ and $\left\|Q\right\|_{\mathrm{F}}$ indicates the Frobenius norm upper bound for weights and gradients respectively. By comparing this bound with Corollary 2,
    • we have $\left\|Q\right\|_{\mathrm{F}}\leq\left\|Q\right\|_{*}\leq\sqrt{r_{G}}\left\|Q\right\|_{\mathrm{F}}$, where $r_{G}$ is the rank of $Q$. Thus when $G_{t}$ are low-rank, or have very imbalanced singular values, $\left\|Q\right\|_{*}$ can be close to $\left\|Q\right\|_{\mathrm{F}}$;
    • we have $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}$ is the max rank of $W_{t}-W_{*}$. Thus, when $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values in a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
    Therefore, we can see in general, ASGO should work well when $G_{t}$ are low-rank and $W_{t}-W_{*}$ are relatively high-rank.
  • Intuitions on ASGO in practical tasks. We argue that low-rank $G_{t}$ and relatively high-rank $W_{t}-W_{*}$ should be common in many practical tasks, revealing the potential of ASGO in applications. As we have discussed in Section 2, gradients are commonly low-rank in deep neural networks, as verified by Zhao et al. (2021); Yang et al. (2023); Cosson et al. (2023). Meanwhile, given the huge success of LoRA (Hu et al., 2022) in fine-tuning foundation models, which naturally asserts a low-rank total update in $W$, there may seem to be a conflict to assume that $W_{t}-W_{*}$ has a high rank. However, it has been observed that $W_{0}-W_{*}$ should be relatively high-rank to obtain a better result, at least in pretraining and some complex fine-tuning tasks for large foundation models (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025). Further exploration of the connection between the rank of weight updates and the performance of algorithms will be an interesting topic.
  • Comparison with Full-Matrix AdaGrad and Shampoo. Shampoo (Gupta et al., 2018) and the full-matrix AdaGrad (Duchi et al., 2011) achieve the following convergence rates under the same settings as Theorem 1.
    Full-Matrix AdaGrad: $\quad\mathcal{O}\left(D_{\mathrm{F}}\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
    Shampoo: $\quad\mathcal{O}\left(\sqrt{r}D_{\mathrm{F}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right)\right).$
    We can check that Theorem 1 indicates a convergence speed that is at least $D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than full-matrix AdaGrad and $\sqrt{r_{G}}D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than Shampoo (proofs are available in Appendix C). This also provides theoretical evidence that single-side preconditioning may better exploit the low-rankness of gradients, showing a faster convergence speed compared to Shampoo-like preconditioning.
@@
The proof is available in Appendix C. We keep the matrix structure of $W_{t}$ and $G_{t}$ throughout our analysis, in contrast with the standard analysis of the convergence of AdaGrad and Shampoo, which is based on vectorizations of $W_{t}$ and $G_{t}$. This is important for proving that ASGO can exploit the structured properties.

### v2

@@ Abstract
Training deep neural networks is a structured optimization problem, because the parameters are naturally represented by matrices and tensors rather than by vectors. Under this structural representation, it has been widely observed that gradients are low-rank and Hessians are approximately block-wise diagonal. These structured properties are crucial for designing efficient optimization algorithms, but are not utilized by many current popular optimizers like Adam. In this paper, we present a novel optimization algorithm ASGO that capitalizes on these properties by employing a preconditioner that is adaptively updated using structured gradients. By fine-grained theoretical analysis, ASGO is proven to achieve superior convergence rates compared to existing structured gradient methods. Based on the convergence theory, we further demonstrate that ASGO can benefit from the low-rank and block-wise diagonal properties. We also discuss practical modifications of ASGO and empirically verify ASGO’s effectiveness on language model tasks.
@@ 1 Introduction
Numerical optimization algorithms, especially those can efficiently train large foundation models (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023, ; Ouyang et al., 2022), play an important role in the modern machine learning field. Among them, adaptive gradient methods like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2014) are popular choices, gaining huge success in training state-of-the-art models in many tasks. These algorithms typically apply a diagonal matrix preconditioner to the gradient $g_{t}$ to update the deep neural network (DNN) parameters $w_{t}$;
@@
This coordinate-wise step size design has been theoretically verified to be effective as it can exploit the sparsity of the gradient vectors $g_{t}$ (Duchi et al., 2011). Also, when the Hessian is well-approximated by a diagonal matrix whose diagonal entries have very different scales, adaptive gradient methods have been proven to be beneficial (Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). While these results seem to be convincing, common DNNs do not necessarily have sparse gradients or Hessians that are well-approximated by ill-conditioned diagonal matrices. Instead, if we take the matrix structure of gradients in neural networks into account, it has been widely observed that these structured gradients are usually low-rank (Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023), and the Hessians are well-approximated by block-wise diagonal matrices (Collobert, 2004; Zhang et al., 2024a, b). Since adaptive gradient methods like Adam, treat the parameters as vectors and ignore the matrix structure of gradients, they are generally unable to exploit these structured properties. This gap makes us ask the following question: How can we properly consider the matrix structures of gradients and exploit their low-rank and block-wise diagonal properties?
One possible answer is provided by Shampoo (Gupta et al., 2018):
$W_{t+1}=W_{t}-\eta_{t}L_{t}^{-\frac{1}{4}}G_{t}R_{t}^{-\frac{1}{4}},\text{ where $W_{t},G_{t}\in\mathbb{R}^{m\times n}$ and $L_{t}\in\mathbb{R}^{m\times m}$, $R_{t}\in\mathbb{R}^{n\times n}$ are full matrices.}$.
@@
The main motivation for such a design is that if we apply vectorization to the update, the Shampoo preconditioner is a single matrix that is the Kronecker product of $L_{t}^{-\frac{1}{4}}$ and $R_{t}^{-\frac{1}{4}}$, which approximates the full-matrix preconditioner for AdaGrad (Duchi et al., 2011). However, the theoretical convergence of Shampoo is worse than that for AdaGrad or even SGD when the dimension is large. Also, Shampoo needs more memory and much heavier computation than adaptive gradient methods because it requires two preconditioners, making it less suitable for training large-scale DNNs.
@@
In this paper, we provide an answer to the aforementioned question by proposing ASGO (Adaptive Structured Gradient Optimization), which significantly improves the convergence guarantees of Shampoo while requiring less memory and computation. In light of the analysis that demonstrates the benefits of adaptive gradient methods (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2020), we use appropriate assumptions to enable a more fine-grained convergence analysis, showing superior convergence results and how ASGO can benefit from the low-rank and block-wise diagonal properties of the problem. We also discuss the connection between ASGO and Muon (Jordan et al., 2024), a structured gradient method based upon the steepest descent algorithm with spectral norm, to conjecture a relation between ASGO and Muon analogous to the relation between AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). Furthermore, we develop an efficient design for query-key attention parameters in transformer models and examine the empirical performance on pretraining transformer model tasks. Our main contributions are summarized as follows.
@@
  • We propose the structured gradient based algorithm ASGO, theoretically analyze its convergence, and show that it converges faster than full-matrix AdaGrad and Shampoo.
@@
  • We further demonstrate that ASGO can effectively exploit the low-rankness of gradients as well as the approximate block-wise diagonal property of Hessians that is typically observed in training DNNS, hence indicating great potential of ASGO for real-world applications.
@@
  • We develop a practical implementation of ASGO with targeted modifications for transformer architectures that offers significant advantages: it eliminates the need for a separate optimizer for 1D parameters in contrast to Muon, while requiring less memory and computational complexity than Shampoo.
@@
  • We empirically validate the effectiveness of ASGO with modifications on language model tasks, demonstrating the algorithm’s great potential in real applications.
@@ 2 Related Work
In the standard optimization literature (i.e., excluding areas such as conic optimization), it is common to consider variables as vectors. However, recently, optimization methods for machine learning that consider variables as matrices have been rapidly gaining attention. Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) consider the matrix or layer structure to help reduce the memory cost of Adam and enable more efficient training. Shampoo (Gupta et al., 2018) and KFAC (Martens and Grosse, 2015) are two pioneering works that approximate, respectively, the full-matrix preconditioner of AdaGrad and the true Fisher matrix (FM), using Kronecker products of smaller matrices, making the memory cost more affordable. Unfortunately, the analysis in (Gupta et al., 2018) shows that the rate of convergence for Shampoo is no better than that for the full-matrix AdaGrad. We note that there is another method, TNT (Ren and Goldfarb, 2021) that is very closely related to Shampoo. TNT was developed as a natural gradient method, approximating the true FM by the covariance of block-wise sampling-based gradients assuming that they are Tensor-Normally distributed. The main difference between TNT and Shampoo, is that TNT uses true FMs and inverses of their Kronecker factors, whereas Shampoo uses empirical FMs and the -1/4 power of their factors. More discussion on recent and concurrent work is presented in Appendix A.
3 Our ASGO Algorithm
3.1 Notation and problem setting
Throughout this paper, we use capital letters such as $W$ to represent matrices and $[W]_{i,j}$ to denote the $(i,j)$-th entry of $W$. For an arbitrary matrix $W\in\mathbb{R}^{m\times n}$, we denote
@@ 3.1 Notation and problem setting
For symmetric square matrices, $A,B\in\mathbb{R}^{m\times m}$, $A\preceq B$ denotes that $B-A$ is positive semidefinite and $A\prec B$ denotes that $B-A$ is positive definite. $\succ$ and $\succeq$ are defined accordingly. We study the following stochastic optimization problem:
@@
where we only have access to a stochastic gradient oracle $\nabla f(W;\xi)$ at $W$.
3.2 ASGO (Algorithm 1)
We propose ASGO (Adaptive Structured Gradient Optimization) in Algorithm 1, which is an algorithm with a single-side preconditioner. Compared to full-matrix AdaGrad, ASGO preserves and utilizes the matrix structure of $W_{t}$ and $G_{t}$, avoiding the huge memory cost for storing its preconditioner. ASGO’s preconditioner consists of a single matrix compared to the two matrices used by Shampoo, leading to the following main update rule:
$W_{t+1}=W_{t}-\eta_{t}V_{t}^{-\frac{1}{2}}G_{t},\text{ where $W_{t}\in\mathbb{R}^{m\times n}$ and $V_{t}\in\mathbb{R}^{m\times m}$ is a full matrix.}$.
ASGO only needs to store one preconditioner matrix and compute its matrix square root and inverse on each iteration, versus two matrices for Shampoo. On the other hand, ASGO’s preconditioner may not be as good an approximation to the full-matrix AdaGrad preconditioner or the empirical Fisher matrix (Gupta et al., 2018; Morwani et al., 2024) as Shampoo’s. However, as we shall see in the following sections, this design can actually better exploit the low-rankness of gradients and the block-wise diagonal nature of Hessians, in terms of achieving better convergence rates.
@@ 3.2 ASGO (Algorithm 1)
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, schedule $\{\eta_{t}\}$ batch size $M\in\mathbb{N}$, and the number of iterations $T$, $\epsilon>0$, ($\epsilon$ should be small, similar to the $\epsilon$ for Adam or AdaGrad.)
@@
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$
@@
7:     $\Lambda_{t}=V_{t}^{\frac{1}{2}}+\epsilon I_{m}$
@@ 4 Nonsmooth Theory
Although DNNs are generally nonconvex globally, convexity may apply locally in some regions. Thus, convex analysis can be helpful for understanding the behavior of DNN training algorithms.
Assumption 1 (Convexity).
$f(\cdot)$ is convex and $W_{*}$ is one of its minimizers.
@@
If we also assume an upper bound for each stochastic sub-gradient such that $\mathbb{E}\left[G_{t}G_{t}^{\top}\right]\preceq Q^{2}$, where $Q\in\mathbb{R}^{m\times m}$ is a positive definite matrix, Theorem 1 also implies
@@
Remark 1.
Note that we treat $D_{\mathrm{op}}$ as a constant here, but it may increase as the iteration number $T$ increases. This can be addressed, for example, by invoking a projection onto a bounded convex set $\mathcal{W}$ with respect to the norm $\left\|\cdot\right\|_{\Lambda_{t}}$ in each iteration, as in AdaGrad (Duchi et al., 2011). $D_{\mathrm{op}}$ would then be bounded by the spectral norm of $\mathcal{W}$. However, since this projection is rarely used in training DNNs, we follow Gupta et al. (2018) and omit it in Algorithm 1. More importantly, this theoretical bound depends on the trace norm of gradients and the spectral norm of weights, showing that the algorithm can make use of the low-rank property of gradients.
One can easily check that this convergence rate for convex nonsmooth problems is $\mathcal{O}(1/\sqrt{T})$, the same as SGD (Zinkevich, 2003) (see below) and AdaGrad (Duchi et al., 2011). Comparison with SGD. The convergence rate for SGD under the assumptions of Corollary 2 is:
@@
where $D_{\mathrm{F}}$ and $\left\|Q\right\|_{\mathrm{F}}$ are the Frobenius norm upper bounds for the weights and gradients, respectively. By comparing this bound with Corollary 2, we have
  • $\left\|Q\right\|_{\mathrm{F}}\leq\left\|Q\right\|_{*}\leq\sqrt{r_{G}}\left\|Q\right\|_{\mathrm{F}}$, where $r_{G}$ is the rank of $Q$. Thus when $G_{t}$ are low-rank, or have very imbalanced singular values, $\left\|Q\right\|_{*}$ can be close to $\left\|Q\right\|_{\mathrm{F}}$;
  • $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}=\text{max rank of}(W_{t}-W_{*})$. Thus, when $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values of a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
Therefore, ASGO should work well when $G_{t}$ are low-rank and $W_{t}-W_{*}$ are relatively high-rank. Intuitions on ASGO in practical tasks. We argue that low-rank $G_{t}$ and relatively high-rank $W_{t}-W_{*}$ should be common in many practical tasks, revealing the potential of ASGO in applications. As we have discussed in Section 2, gradients are commonly low-rank in DNNs, as verified by Zhao et al. (2021); Yang et al. (2023); Cosson et al. (2023). Meanwhile, given the huge success of LoRA (Hu et al., 2022) in fine-tuning foundation models, which naturally asserts a low-rank total update in $W$, there may seem to be a conflict to assume that $W_{t}-W_{*}$ has a high rank. However, it has been observed that $W_{0}-W_{*}$ should be relatively high-rank to obtain a better result, at least in pretraining and some complex fine-tuning tasks for large foundation models (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025). Further exploration of the connection between the rank of weight updates and the performance of algorithms is an interesting topic for future research. Comparison with Full-Matrix AdaGrad and Shampoo. Shampoo (Gupta et al., 2018) and the full-matrix AdaGrad (Duchi et al., 2011) achieve the following convergence rates under the same settings as Theorem 1.
[…]
We can check that Theorem 1 indicates a convergence speed that is at least $D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than full-matrix AdaGrad and $\sqrt{r_{G}}D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than Shampoo; (see Appendix E for proofs). This also provides theoretical evidence that single-side preconditioning may be better at exploiting low-rankness of gradients, yielding a faster convergence speed compared to Shampoo-like preconditioning.
@@
Proofs of all of the results in this section are given in Appendix E. We keep the matrix structure of $W_{t}$ and $G_{t}$ throughout our analysis, in contrast to standard analyses of the convergence of AdaGrad and Shampoo, which are based on vectorizations of $W_{t}$ and $G_{t}$. This is important for proving that ASGO can exploit the structured properties.